The question is fairly simple. Finding the right order when "of" is concerned throws me for a loop. I assume the common denominator is the right guess but here goes. What is 1/3 of 6/5 of -15/4. Thanks
Answer by Alan3354(69443)

What is 1/3 of 6/5 of -15/4.
============
6/5 of -15/4 means (6/5)*(-15/4) = -90/20 = -9/2
----
1/3 of that = -3/2